# Guido Franchetti1 and Derek Harland2

### v1

@@ 2.3 The moduli space metric
A tangent vector to the framed moduli space at $(A,\Phi)$ is described by an $\mathfrak{su}(2)$-valued 1-form $a$ and 0-form $\phi$ solving the linearised Bogomolny equations,
$d^{A}\phi+[a,\Phi]=\ast d^{A}a,$ (20)
such that $a+\phi\,d\theta$ extends to a smooth circle-invariant section of $\Lambda^{1}(S^{4})\otimes\operatorname{End}(V)$. Given any $\mathfrak{su}(2)$-valued function $\chi$ that extends to a circle-invariant section over $S^{4}$, $(a,\phi)=(d^{A}\chi,[\Phi,\chi])$ is a solution of (20) representing an infinitesimal gauge transformation. A short calculation shows that if $g_{s}$ is a family of gauge transformations such that $g_{0}=0$, $g_{s}^{-1}\partial_{s}g_{s}|_{s=0}=\chi$ then
$\left.\frac{d}{ds}\right|_{s=0}CS[A,g_{s}^{-1}Ag_{s}+g_{s}^{-1}dg_{s}]=2\int_{H^{3}}\operatorname{Tr}(F^{A}\wedge d^{A}\chi).$ (21)
Thus we define the space $G^{u}_{A,\Phi}$ of infinitessimal unframed gauge transformations to be the set of all pairs $(d^{A}\chi,[\Phi,\chi])$ such that $\chi:H^{3}\to\mathfrak{su}(2)$ extends smoothly to a circle-invariant section of $\operatorname{End}(V)\to S^{4}$. We define the space $G^{f}_{A,\Phi}$ of infinitessimal framed gauge transformations to be the subspace such that $\int_{H^{3}}\operatorname{Tr}(F^{A}\wedge d^{A}\chi)=0$. These are both subspaces of the space $T_{A,\Phi}C_{n,p}$ of pairs $(a,\phi)$ that extend smoothly to $S^{4}$ in a circle-invariant way and satisfy equation (20). The tangent space to the framed and unframed moduli spaces are the quotient spaces
$T_{A,\Phi}M^{f}_{n,p}=T_{A,\Phi}C_{n,p}/G^{f}_{A,\Phi},\quad T_{A,\Phi}M^{u}_{n,p}=T_{A,\Phi}C_{n,p}/G^{u}_{A,\Phi}.$ (22)
An $L^{2}$ metric on the moduli space is a metric obtained by integrating $X=(a,\phi)$:
@@
In order to make this well-defined, we must choose a particular representative $(a,\phi)$ of a gauge equivalence class; in other words, we must fix a gauge. Unfortunately, any real representative of a non-zero tangent vector results in a divergent integral. This is because unframed hyperbolic monopoles have the holographic property of being determined by the abelian connection induced at infinity [7]. Any representative $a$ of a non-zero tangent vector must be non-zero when restricted to the line bundles $L^{\pm}$ over $S^{2}_{\infty}$. It follows that $a^{\prime}=-\operatorname{Tr}(a\Phi)/\sqrt{-\operatorname{Tr}(\Phi^{2})}$ gives a non-vanishing 1-form on any geodesic 2-sphere of sufficiently large radius. The metric $dr^{2}+\sinh^{2}(r)g_{S^{2}}$ of $H^{3}$ is such that the 1-form $a^{\prime}$ on $S^{2}$ satisfies $\ast a^{\prime}=\ast_{S^{2}}a^{\prime}\wedge dr$. Thus there exist positive constants $R,C$ such that
$g(X,X)\geq\frac{1}{2}\int_{r\geq R}a^{\prime}\wedge\ast_{S^{2}}a^{\prime}\wedge dr\geq\int_{R}^{\infty}C\,dr>\infty.$ (24)
@@
One way to avoid this divergence is to choose a non-real representative of a real tangent vector. Using an $\mathfrak{sl}(2,\mathbb{C})$-valued gauge transformation, we can choose a representative $(a,\phi)$ that is not antihermitian, but is such that $(a+a^{\dagger},\phi+\phi^{\dagger})$ is pure gauge. The advantage of doing so is that it opens up the possibility of $a|_{S^{2}_{\infty}}$ being a holomorphic (or antiholomorphic) 1-form. A 1-form $a_{\infty}$ that is of type $(1,0)$ or $(0,1)$ with respect to the complex structure of $S^{2}_{\infty}$ satisfies $a_{\infty}\wedge\ast_{S^{2}}a_{\infty}=0$. Thus the divergence of the $L^{2}$ metric could be avoided if the restriction of $a$ to $S^{2}_{\infty}$ were holomorphic or antiholomorphic.
@@
It differs from the more familiar harmonic gauge-fixing condition used in [7, 30] by the term ±$\pm 2i\phi_{\pm}$. Equation (25) is not one but two gauge-fixing conditions corresponding to the choice of sign. The two representatives $(a_{\pm},\phi_{\pm})$ of a real tangent vector to the moduli space will differ by a gauge transformation. We are free to choose $(a_{+},\phi_{+})=(-a_{-}^{\dagger},-\phi_{-}^{\dagger})$, as this choice clearly solves the “+” gauge fixing condition if $(a_{-},\phi_{-})$ solves the “--” condition, and, if $(a_{-},\phi_{-})$ represents a real tangent vector, it differs from $(-a_{+}^{\dagger},-\phi_{+}^{\dagger})$ by a gauge transformation — compare with (29) below.
@@
Let $(A,\Phi)$ be a hyperbolic monopole with half-integer mass and let $(a,\phi)$ be a solution of the linearised Bogomolny equation (20) that represents a tangent vector to the moduli space. Then there exists an unframed $\mathfrak{sl}(2,\mathbb{C})$ gauge transformation $\chi$ such that $(a_{-},\phi_{-})=(a+d^{A}\chi,\phi+[\Phi,\chi])$ satisfies the gauge-fixing condition (25). This $\chi$ has the following properties:
@@
In order to apply this theorem to the framed moduli space, we need the gauge transformation $\chi$ to be framed. But it is easy to modify $\chi$ so that it is framed. It is straightforward to check that if $(a_{-},\phi_{-})$ solves the gauge-fixing condition (25) then so does $(a_{-}-c\,d^{A}\Phi,\phi_{-})$ for any constant $c$. So replacing $\chi$ in Theorem 3 with $\tilde{\chi}=\chi-c\Phi$ and setting
@@
Let $(A,\Phi)$ and $(a,\phi)$ be as in Theorem 3. Then there exists a unique framed gauge transformation $\chi$ such that $(a_{-},\phi_{-})=(a+d^{A}\chi,\phi+[\Phi,\chi])$ satisfies the gauge-fixing condition (25). This $\chi$ has the two properties listed in Theorem 3.
We will show that $(a_{-}-c\,d^{A}\Phi,\phi_{-})$ has the two properties listed in Theorem 3 at the end of section 3. The uniqueness part of this Theorem will be proved in section 4. Turning to the unframed moduli space, we have that:
@@
Let $(A,\Phi)$ and $(a,\phi)$ be as in Theorem 3. Then there exists a unique unframed gauge transformation $\chi$ such that $(a_{-},\phi_{-})=(a+d^{A}\chi,\phi+[\Phi,\chi])$ satisfies the gauge-fixing condition (25) and $\int_{H^{3}}\operatorname{Tr}(a_{-}\wedge F^{A})=0$. This $\chi$ has the two properties listed in Theorem 3.
@@
Theorem 4 yields two symmetric bilinear forms $g_{\pm}$ on the tangent bundle of the framed monopole moduli space. They are defined by
@@
in which $(a_{-},\phi_{-})$ is the solution of the gauge fixing condition (25) determined in Theorem 4, and $(a_{+},\phi_{+})=(-a_{-}^{\dagger},-\phi_{-}^{\dagger})$. By construction, the bilinear forms $g_{\pm}$ are also related by conjugation: in other words,
@@
for any real tangent vector $X$.
We have been careful to avoid calling $g_{\pm}$ metrics because we do not know whether they always real-valued, and if so, whether they are positive definite. They are real if and only if $g_{+}=g_{-}$, so we can investigate reality by calculating their difference. To do this, it is convenient to write
@@
More can be said in the case of inversion-symmetric monopoles. Recall that inversion about a point in $H^{3}$ is the unique orientation-reversing isometry $I:H^{3}\to H^{3}$ whose fixed set is a point and which squares to the identity. A monopole is called inversion-symmetric if
@@ 3 Existence for gauge fixing condition
The purpose of this section is to prove Theorem 3, in other words, to show that the gauge-fixing condition (25) has a unique solution. We begin by reformulating it in a 4-dimensional way. Let $g_{H}=d\theta^{2}+g_{H^{3}}$ be the metric on $S^{1}\times H^{3}$ and let $\ast_{H}$ be the associated Hodge star. Let $B=\Phi\,d\theta+A$ be the circle-invariant connection on $S^{4}$ and let $b=\phi\,d\theta+a$ be the deformation of $B$ associated to $(a,\phi)$. Then, for the choice of orientation $d\theta\wedge\operatorname{Vol}_{H^{3}}$, (25) is equivalent to $\delta^{\pm}b=0$, where
$\delta^{\pm}:\Omega^{1}(\mathfrak{su}(2))\to\Omega^{0}(\mathfrak{su}(2)),\quad\delta^{\pm}b=e^{\mp 4i\theta}\ast_{H}d^{B}\ast_{H}e^{\pm 2i\theta}b$ (40)
Here the factor of $e^{\mp 4i\theta}$ is included for later notational convenience.
We have already seen in (7) that the metric $g_{H}$ on $S^{1}\times H^{3}$ is conformally equivalent to spherical metric $g_{S}=dy_{\mu}dy_{\mu}$ on $S^{4}\setminus S^{2}$. To begin analysing equation (40) we employ a second conformal equivalence. Introducing $r_{2}^{2}=y_{3}^{2}+y_{4}^{2}+y_{5}^{2}$, we note that
$\frac{1}{r_{2}^{2}}dy_{\mu}dy_{\mu}=\frac{1}{r_{2}^{2}}(dr_{2}^{2}+dy_{1}^{2}+dy_{2}^{2})+\frac{1}{r_{2}^{2}}(dy_{3}^{3}+dy_{4}^{2}+dy_{5}^{2})$ (41)
is the product metric on $H^{3}\times S^{2}$, with $S^{2}$ consisting of unit vectors $(y_{3},y_{4},y_{5})/r_{2}$. The $S^{4}$ constraint $y_{\mu}y_{\mu}=1$ is equivalent to $r_{2}^{2}+y_{1}^{2}+y_{2}^{2}=1$, which defines a copy of $H^{2}$ inside $H^{3}$. So the metric $g_{S}$ on $S^{4}\setminus S^{2}$ is conformal to the product metric $g_{K}$ on $H^{2}\times S^{2}$. It is convenient to work in the disc model of $H^{2}$, so we introduce a stereographic coordinate $w\in\mathbb{C}$ satisfying $|w|<1$, such that
@@
Then the argument of $w$ equals $\theta$ and the circle action fixes the $S^{2}$ given by $w=0$. The product metric on $H^{2}\times S^{2}$ obtained from (41) is
@@
with $z\in\mathbb{C}$ being a stereographic coordinate on $S^{2}$. Since $r_{2}=(1-|w|^{2})/(1+|w|^{2})$ and $r_{1}=2|w|/(1+|w|^{2})$, the metric $g_{H}$ of equation (7) and the spherical metric $g_{S}$ are related to $g_{K}$ by
@@
Moreover, the natural orientations of $S^{1}\times H^{3}$ and $H^{2}\times S^{2}$ agree.
Like $g_{S}$, the metric $g_{K}$ extends smoothly across the $S^{2}\subset S^{4}$ fixed by the circle action. We will use the metric $g_{K}$ to show that the gauge-fixing condition $\delta^{-}b=0$ admits a solution in a neighbourhood of this $S^{2}$, before showing that a solution exists on all of $S^{4}$. Solving the alternative condition $\delta^{+}b=0$ is equivalent, and we focus on $\delta^{-}$ for notational convenience. The advantage of working with $g_{K}$, rather than directly with the round metric $g_{S}$, is that $g_{K}$ is Kähler, with Kähler form

### v2

@@ 2.3 The moduli space metric
An $L^{2}$ metric on the moduli space is a metric obtained by integrating the pointwise inner product of a tangent vector $X=(a,\phi)$ with itself:
@@
In order to make this well-defined, we must choose a particular representative $(a,\phi)$ of a gauge equivalence class; in other words, we must fix a gauge. Unfortunately, any real representative of a non-zero tangent vector results in a divergent integral. This is because unframed hyperbolic monopoles have the holographic property of being determined by the abelian connection induced at infinity [7]. Any representative $a$ of a non-zero tangent vector must be non-zero when restricted to the line bundles $L^{\pm}$ over $S^{2}_{\infty}$. It follows that $a^{\prime}=-\operatorname{Tr}(a\Phi)/\sqrt{-\operatorname{Tr}(\Phi^{2})}$ gives a non-vanishing 1-form on any geodesic 2-sphere of sufficiently large radius. The metric $dr^{2}+\sinh^{2}(r)g_{S^{2}}$ of $H^{3}$ is such that the 1-form $a^{\prime}$ on $S^{2}$ satisfies $\ast a^{\prime}=\ast_{S^{2}}a^{\prime}\wedge dr$. Hence there exist positive constants $R,C$ such that
$g(X,X)\geq\frac{1}{2}\int_{r\geq R}a^{\prime}\wedge\ast_{S^{2}}a^{\prime}\wedge dr\geq\int_{R}^{\infty}C\,dr=\infty.$ (48)
@@
One way to avoid this divergence is to choose a non-real representative of a real tangent vector. Using an $\mathfrak{sl}(2,\mathbb{C})$-valued gauge transformation, we can choose a representative $(a,\phi)$ that is not antihermitian, but is such that $(a+a^{\dagger},\phi+\phi^{\dagger})$ is pure gauge. The advantage of doing so is that it opens up the possibility of $a|_{S^{2}_{\infty}}$ being a holomorphic (or antiholomorphic) 1-form. A 1-form $a_{\infty}$ that is of type $(1,0)$ or $(0,1)$ with respect to the complex structure of $S^{2}_{\infty}$ satisfies $a_{\infty}\wedge\ast_{S^{2}}a_{\infty}=0$. Thus the divergence of the $L^{2}$ metric could be avoided if the restriction of $a$ to $S^{2}_{\infty}$ were holomorphic or antiholomorphic. Note that this argument only applies to the complex-bilinear form $g$ defined in (47); it does not apply to the hermitian metric $\int\operatorname{Tr}(a\wedge\ast\bar{a}+\phi\ast\bar{\phi})$ because $a_{\infty}\wedge\ast_{S^{2}}\bar{a}_{\infty}=0$ only when $a_{\infty}=0$.
@@
It differs from the more familiar harmonic gauge-fixing condition used in [7, 32] by the term $\pm 2i\phi_{\pm}$. Equation (49) is not one but two gauge-fixing conditions corresponding to the choice of sign. The two representatives $(a_{\pm},\phi_{\pm})$ of a real tangent vector to the moduli space will differ by a gauge transformation. We are free to choose $(a_{+},\phi_{+})=(-a_{-}^{\dagger},-\phi_{-}^{\dagger})$, as this choice clearly solves the “+” gauge fixing condition if $(a_{-},\phi_{-})$ solves the “-” condition, and, if $(a_{-},\phi_{-})$ represents a real tangent vector, it differs from $(-a_{+}^{\dagger},-\phi_{+}^{\dagger})$ by a gauge transformation — compare with (53) below.
@@
Let $(A,\Phi)$ be a hyperbolic monopole with half-integer mass and let $(a,\phi)$ be a solution of the linearised Bogomolny equation (43) that represents a tangent vector to the moduli space. Then there exists an infinitesimal unframed $\mathfrak{sl}(2,\mathbb{C})$ gauge transformation $\chi$ such that $(a_{-},\phi_{-})=(a+d^{A}\chi,\phi+[\Phi,\chi])$ satisfies the gauge-fixing condition (49). This $\chi$ has the following properties:
@@
In order to apply this theorem to the framed moduli space, we need the gauge transformation $\chi$ to be framed. But it is easy to modify $\chi$ so that it is framed. If $(a_{-},\phi_{-})$ solves the gauge-fixing condition (49) then so does $(a_{-}-c\,d^{A}\Phi,\phi_{-})$ for any constant $c$. So replacing $\chi$ in Theorem 12 with $\tilde{\chi}=\chi-c\Phi$ and setting
@@
Let $(A,\Phi)$ and $(a,\phi)$ be as in Theorem 12. Then there exists a unique infinitesimal framed gauge transformation $\chi$ such that $(a_{-},\phi_{-})=(a+d^{A}\chi,\phi+[\Phi,\chi])$ satisfies the gauge-fixing condition (49). This $\chi$ has the two properties listed in Theorem 12.
We will show that $(a_{-}-c\,d^{A}\Phi,\phi_{-})$ has the two properties listed in Theorem 12 at the end of section 3. The uniqueness part of this Theorem will be proved in Proposition 28 in section 4. Turning to the unframed moduli space, we have that:
@@
Let $(A,\Phi)$ and $(a,\phi)$ be as in Theorem 12. Then there exists a unique infinitesimal unframed gauge transformation $\chi$ such that $(a_{-},\phi_{-})=(a+d^{A}\chi,\phi+[\Phi,\chi])$ satisfies the gauge-fixing condition (49) and $\int_{H^{3}}\operatorname{Tr}(a_{-}\wedge F^{A})=0$. This $\chi$ has the two properties listed in Theorem 12.
@@
Theorem 13 ensures that calculating the bilinear form (47) for $(a,\phi)$ solutions of the gauge fixing condition (49) yields two symmetric bilinear forms $g_{\pm}$ on the tangent bundle of the framed monopole moduli space. They are given by
@@
for $(a_{-},\phi_{-})$ the solution of the gauge fixing condition (49) determined in Theorem 13, and $(a_{+},\phi_{+})=(-a_{-}^{\dagger},-\phi_{-}^{\dagger})$. By construction, the bilinear forms $g_{\pm}$ are also related by conjugation: in other words,
@@
for any real tangent vector $X$. We recall that we call a tangent vector $X=(a,\phi)$ real if $a+a^{\dagger}$, $\phi+\phi^{\dagger}$ are pure gauge.
We have been careful to avoid calling $g_{\pm}$ metrics because we do not know whether they are always real-valued, and if so, whether they are positive definite. They are real if and only if $g_{+}=g_{-}$, so we can investigate reality by calculating their difference. To do this, it is convenient to write
@@
More can be said in the case of inversion-symmetric monopoles. Recall that inversion about a point in $H^{3}$ is the unique orientation-reversing isometry $I:H^{3}\to H^{3}$ whose fixed set is that point and which squares to the identity. A monopole is called inversion-symmetric if
@@
Remark 16.
The reader may wonder whether the constant $\pm 2i$ in the gauge fixing condition (54) could be replaced by another constant. The choice $\pm 2i$ is special because it is linked to supersymmetry and leads to special geometry, which will be described in Section 4. Other choices close to $\pm 2i$ might lead to convergent metrics, but those metrics would not be expected to have any special geometric structure. ∎
Remark 17.
The reader may recall that the Coulomb gauge fixing condition $\ast d^{A}\ast a+[\Phi,\phi]=0$ used to define the euclidean monopole metric has a natural interpretation: it says that $(a,\phi)$ is orthogonal to infinitesimal gauge transformations. We do not know of a comparable interpretation for the gauge fixing condition (49). ∎
@@ 3 Existence for gauge fixing condition
3.1 Proof strategy
The purpose of this section is to prove the existence part of Theorems 12 and 13. In other words, to show that the gauge-fixing condition (49) has a solution satisfying the required properties.
We begin by reformulating (49) in a 4-dimensional way. Let $g_{H}=d\theta^{2}+g_{H^{3}}$ be the metric on $S^{1}\times H^{3}$ and let $\ast_{H}$ be the associated Hodge star. Let $B=\Phi\,d\theta+A$ be a circle-invariant connection on $S^{1}\times H^{3}$ and let $b=\phi\,d\theta+a$ be the deformation of $B$ associated to $(a,\phi)$. Recall that both $B$ and $b$ extend to $S^{4}$. Then, for the choice of orientation $d\theta\wedge\operatorname{Vol}_{H^{3}}$, the gauge fixing condition (49) is equivalent to
$\delta^{\pm}b=0,$ (64)
where
$\delta^{\pm}:\Omega^{1}(\mathfrak{su}(2))\to\Omega^{0}(\mathfrak{su}(2)),\quad\delta^{\pm}b=e^{\mp 4i\theta}\ast_{H}d^{B}\ast_{H}e^{\pm 2i\theta}b.$ (65)
Here the factor of $e^{\mp 4i\theta}$ is included for later notational convenience. We need to find an infinitesimal gauge transformation $\chi$ on $S^{4}$ such that $b-d^{B}\chi$ solves
$\delta^{\pm}(b-d^{B}\chi)=0$ (66)
on $S^{4}\setminus S^{2}$.
The way we solve (66) is in two steps. First we find $\chi_{1}$ such that $\delta^{\pm}(b-d^{B}\chi_{1})$ equals 0 on $S^{2}\subset S^{4}$ and decays rapidly as one approaches $S^{2}$. We do this in section 3.2 and our argument relies on the fact that $H^{3}\times S^{1}$ is conformally related to the Kähler manifold $H^{2}\times S^{2}$. The second step is to use elliptic theory to find $\chi_{2}$ solving $\delta^{\pm}(b-d^{B}\chi_{1}-d^{B}\chi_{2})=0$ on $S^{4}$. We do this in section 3.3, where we also show that the solution of the gauge fixing condition that we have constructed satisfies the properties listed in Theorems 12, 13. In the reminder of this section we describe the aforementioned conformal equivalence $H^{3}\times S^{1}\simeq H^{2}\times S^{2}$.
We have already seen in (13) that the metric $g_{H}$ on $S^{1}\times H^{3}$ is conformally equivalent to metric $g_{S}$ on $S^{4}\setminus S^{2}$. Recall that we model $S^{4}$ as the set of vectors $y\in\mathbb{R}^{5}$ satisfying $y_{\mu}y_{\mu}=1$. The metric $g_{S}$ is the restriction of the euclidean metric $dy_{\mu}dy_{\mu}$ to $S^{4}$. Introducing $r_{2}^{2}=y_{3}^{2}+y_{4}^{2}+y_{5}^{2}$, we note that
$\frac{1}{r_{2}^{2}}dy_{\mu}dy_{\mu}=\frac{1}{r_{2}^{2}}(dr_{2}^{2}+dy_{1}^{2}+dy_{2}^{2})+\frac{1}{r_{2}^{2}}(dy_{3}^{2}+dy_{4}^{2}+dy_{5}^{2})$ (67)
is the product metric on $H^{3}\times S^{2}$, with $r_{2}>0,y_{1},y_{2}$ being coordinates on $H^{3}$ and $S^{2}$ consisting of unit vectors $(y_{3},y_{4},y_{5})/r_{2}$. The $S^{4}$ constraint $y_{\mu}y_{\mu}=1$ is equivalent to $r_{2}^{2}+y_{1}^{2}+y_{2}^{2}=1$, which determines a submanifold of $H^{3}$ that is isometric to $H^{2}$. So $S^{4}\setminus S^{1}$ is conformal to $H^{2}\times S^{2}$, with $S^{1}\subset S^{4}$ defined by $r_{2}=0$.
We let $g_{K}$ denote the metric on $H^{2}\times S^{2}$. It is convenient to write it using the disc model of $H^{2}$. To do so we introduce a stereographic coordinate $w\in\mathbb{C}$ satisfying $|w|<1$, such that
@@ 3.1 Proof strategy
The metric (67) on $H^{2}\times S^{2}$ then becomes
@@
with $z\in\mathbb{C}$ being a stereographic coordinate on $S^{2}$. It is important to note that $g_{K}$ is Kähler, with Kähler form
We now have three conformally equivalent metrics $g_{H},g_{S},g_{K}$. Their common domain of definition is $S^{4}\setminus(S^{2}\sqcup S^{1})\simeq S^{1}\times(H^{3}\setminus\{0\})\simeq(H^{2}\setminus\{0\})\times S^{2}$, with 0 denoting points in the interior of $H^{3}$ and $H^{2}$. From (13) and (67) we have $g_{H}=r_{1}^{-2}g_{S}$ and $g_{K}=r_{2}^{-2}g_{S}$. Since $r_{2}=(1-|w|^{2})/(1+|w|^{2})$ and $r_{1}^{2}=1-r_{2}^{2}$, the metric $g_{H}$ of equation (13) and the spherical metric $g_{S}$ are related to $g_{K}$ by
@@
We note that the argument of the coordinate $w$ on $H^{2}$ equals the coordinate $\theta$ on $S^{1}$: If we define $s:=-\ln|w|>0$ then we find from (71) and (69) that
$g_{H}=d\theta^{2}+ds^{2}+\sinh^{2}(s)\frac{2}{(1+|z|^{2})^{2}}(dz\,d\bar{z}+d\bar{z}\,dz).$ (72)
So $s$ is a geodesic radial coordinate on $H^{3}$, i.e. it measures the geodesic distance from the origin. Therefore we can write $w=e^{-s+i\theta}$, with $s$ being a radial coordinate on $H^{3}$ and $\theta$ being the coordinate on $S^{1}$.
3.2 Solution on $S^{2}$
The aim of this section is to find an infinitesimal gauge transformation $\chi$ such that
$(b-d^{B}\chi)^{0,1}=w^{2p}\sigma,$ (73)
for $\sigma\in\Gamma(\Lambda^{0,1}(S^{2})\otimes\operatorname{End}_{0}(V))$, and such that
$\frac{1}{w^{3}}\delta^{-}(b-d^{B}\chi)$ (74)
is a smooth function on $S^{4}$. In particular $(b-d^{B}\chi)^{0,1}$ and $\delta^{-}(b-d^{B}\chi)$ vanish on the 2-sphere fixed by the circle action and corresponding to $w=0$.